Dear colleagues,
Here is a summary of today's activites at CSBF.
- The MXR problem did not disappear during the overnight run. Jim came in the morning and confirmed which cable causes the time out error. The cable lengths between concentrators are checked and sent to Nick. A newly made cable (10 pin flat twisted cable) by Isaac&Nick has been sent via FedEx priority.
- SCD heat pipe fitting was given to Fritz. He has spent time to figure out how to install them. Dave and Fritz talked with Scott on his trip schedule for the installation.
- -X side, sparsification boards stack and master trigger board, has cables cleaned up. So besides SFCs which will be done after the SSDs delivered and integrated, all cablings of instrument are cleaned. I uploaded pictures of cabling on the black board under CSBF integration.
- The disconnected Chercam HSK MUX connector replaced and saw most of channels in blue colors except 8 HV values. Asked Laurent if they are OK or not.
- TCD Power hsk mapping conversion modified to have correct display. Still there is a saturation problem with the 12 V monitoring channel. Asking Isaac for their pin mapping for clarification. Other channels are OK now.
- Second boot source problem with the HSK board solved by Prashant with Larry's advice.
- Debugging of the reduced hsk data being worked by Prashant.
- SangEun has been working on the SIP interface coding.
- Cleaning up of integration data inside 3 client PCs has started. Going to have enough back ups of CREAM4 integration data in 3 PCs. Those data taken at UMD seems too big (maybe ~ 20 GB) to be transferred in one or two day. If they don't have to be copied later, we can plan to bring those data by an USB external disk when we return to UMD. Please let us know whether you need those data before that or not.
Here are things to do.
- Reduced hsk test with the Iridium connection from CDM setup.
- Integration with the CDM.
- Installation of SSDs into SFC1/2 if delivered.
- Power cable to CDM installation if delivered.
- CD housekeeping conversion factor in the housekeeping table.
- TCD Power +12V mapping debugging.
- Update of manuals including CPT.
- TCD MXR re-cabling when delivered.
- TCD S3 tuning.
- CD integration from Thursday.
- Daily data copying to the cdps4 until we can have 2 Ethernet cards and make it working.
Please let us know if you have any questions/comments/suggestions.
Cheers, MooHyun.
MooHyun, July 16, 2008



